OpenStudy (hari5719):

5x^2 +x-7=0 giving each solution correct to 2 decimal places

OpenStudy (hari5719):

@tayromo

OpenStudy (hari5719):

@abb0t

OpenStudy (thomas5267):

Use quadratic formula.

OpenStudy (hari5719):

i want the answer not the way sorry

OpenStudy (cp9454):

\[\frac{ -1\pm \sqrt{141} }{ 10 }\]

OpenStudy (cp9454):

-1.28 and 1.08

OpenStudy (hari5719):

and the working

OpenStudy (hari5719):

@tayromo can u work it out

OpenStudy (anonymous):

they are giving you the answers right above you. sorry i was working on something else

OpenStudy (hari5719):

i want the working out @tayromo @cp9454

OpenStudy (anonymous):

You have an equation in the quadratic form.

OpenStudy (anonymous):

\[ax^2+bx+c=0\] \[5x^2+x-7=0\]

OpenStudy (anonymous):

So you can use the quadratic equation.

OpenStudy (anonymous):

Quadratic equation: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Identify your a, b and c then plug and solve.

OpenStudy (hari5719):

can u work it out pls @Johnbc

OpenStudy (hari5719):

@yanee @Squirrels

OpenStudy (hari5719):

@AmericaS

OpenStudy (hari5719):

can someone work it out and tell the answer plssssss

OpenStudy (americas):

Um, math isn't my best subject. :( Sorry.

OpenStudy (hari5719):

its ok

OpenStudy (hari5719):

@cp9454 can u work it and explain

OpenStudy (anonymous):

it has been worked out Hari. They gave you the equation and the answers. They are just in separate boxes

OpenStudy (hari5719):

lol not the working out

OpenStudy (anonymous):

All you have to do is plug the values in and solve.

OpenStudy (hari5719):

work it out plssssssssss

OpenStudy (hari5719):

which is a b c

OpenStudy (anonymous):

They correspond to the coefficients of the quadratic form. So for your case compare the quadratic form with your quadratic and you will see that a = 5 b = 1 c = -7

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

@hari5789 read the rules.

OpenStudy (anonymous):

\[5x^2+x-7=0\] quadratic formulea \[x=-b \pm \sqrt{b^2-4ac} / 2a\] a=5, b=1, c=-7 \[x=-1 \pm \sqrt{1^2-4(5)(-7)}/2(5)\] \[x=-1 \pm \sqrt{141}/10\] \[x=-1-\sqrt{141}/10=-1.29\] \[x=-1+\sqrt{141}/10=1.09\] so the answer is x=-1.29 and x=1.09